OpenStudy (anonymous):

simplify, leaving your answer in surd form. 84√12 / 42√2?

OpenStudy (anonymous):

divided

OpenStudy (mertsj):

\[\frac{84\sqrt{12}}{42\sqrt{2}}=\frac{84}{42}\times \sqrt{\frac{12}{2}}=2\sqrt{6}\]

OpenStudy (anonymous):

can you help me on a few more?

OpenStudy (mertsj):

\[\frac{\sqrt{3}}{2}\]

OpenStudy (mertsj):

\[8\sqrt{90}-4\sqrt{40}-4\sqrt{10}=8\sqrt{9(10)}-4\sqrt{4(10)}-4\sqrt{10}=24\sqrt{10}-8\sqrt{10}-4\sqrt{10}\]

OpenStudy (mertsj):

\[24\sqrt{10}-8\sqrt{10}-4\sqrt{10}=12\sqrt{10}\]

OpenStudy (anonymous):

simplify: 4√50 + 7√32 - 7√2 ?

OpenStudy (anonymous):

4√8 + 6√18 + 7√2 ? simplify?